I need help on how to do find the answer to this problem (5-3i)(1-2i)
Answer by stanbon(57379)

(5-3i)(1-2i)
---
= 5(1-2i)-3i(1-2i)
--------
= 5-10i-3i+6(i^2)
---
= 5-13i+6(-1)
---
= 5-6-13i
---
= -1-13i
=========
Cheers,
Stan H.
=========